OpenStudy (anonymous):

A bug walking on the circular rim of a flower pot completes one lap around the planter in 12.3 s. If the radius of the planter is 11.5 cm, how fast was it traveling? Answer in units of cm/s How fast in m/s? Answer in units of m/s

OpenStudy (anonymous):

Think of the rim of the flower pot as being a circle. Imagine unraveling that circle an turning it into a straight line. That's how far the bug travelled. The distance it traveled is therefore equal to the circumference of the circle. Do \[d = 2 \pi r\] where r is the radius of the pot. Now, to find out how fast it traveled, recall the units for velocity. \[v = m/s\] velocity is meters per second. Therefore, to find out how fast it traveled, divide the distance it went, by the time it took. Do \[v = d/t\] Using 11.5cm, as given in the problem, will give you the answer in cm/s. Divide that answer by 100 to get the answer in m/s.